A silver nitrate solution contains 14.77 g of primary-standard AgNO3 in 1.00 L. What volume of this solution will be needed to react with 0.1957g of NaCl.
1
Expert's answer
2020-10-21T12:52:05-0400

The balanced reaction equation is:

AgNO3 + NaCl "\rightarrow" NaNO3 + AgCl"\downarrow" .

As one can see, 1 mol of silver nitrate reacts with 1 mol of sodium chloride. The number of the moles of sodium chloride can be calculated from the ratio of its mass to its molar mass (58.44 g/mol):

"n(NaCl) = \frac{m(NaCl)}{M(NaCl)} = \frac{0.1957 \text{ g}}{58.44 \text{ g/mol}} = 0.003349" mol.

Therefore, 0.003349 mol of silver nitrate will be needed to react with 0.1957g of NaCl. Let's calculate the mass of silver nitrate that contains 0.003349 mol (molar mass of AgNO3 is 169.87 g/mol):

"m(AgNO_3) = n\cdot M = 0.003349\cdot169.87 = 0.5688" g.

If a silver nitrate solution contains 14.77 g of AgNO3 in 1L, then the volume needed is:

"V = \frac{0.5688 \text{ g }\cdot1 \text{L}}{14.77\text{ g}} = 0.03851" L, or 38.5 mL.

Answer: 38.5 mL of the silver nitrate solution (14.77 g/L) will be needed to react with 0.1957 g of NaCl.